# SySTeC: A Symmetric Sparse Tensor Compiler

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Mathematical softwareccs: Software and its engineering Compilersccs: Computing methodologies Symbolic and algebraic manipulation
@@ 3.2.1. Visible Output Symmetry
For example, let us consider the first block of the symmetrized TTM kernel given in Listing 1 that performs the assignments using coordinates of $A$ in the canonical triangle that are not on a diagonal. We reorder the assignments to make the pattern from output symmetry more obvious in Listing 2. Swapping the second and third indices in the output tensor on the left-hand side lends an equivalent right-hand side for each expression. As depicted in Listing 3, we can exploit the output symmetry by only writing to the canonical triangle of the output tensor (i.e. if we index C as C[i, j, l], then only where j <= l), which reduces the number of computations that are done by $\frac{1}{2}$. Then, we can copy the values from the canonical triangles to the other triangles of the output tensor in a separate loop nest (lines 7-9 of Listing 3), thus completing the remaining assignments.
@@ 9 C[i, j, l] = C[i, l, j]
In general, if $n$ indices are in the same part of a partition representing the visible symmetry of the output tensor, then we can reduce the number of assignment operations by $\frac{1}{n!}$.
@@ 4 A = A[i, j]
If $n$ indices are in the same part of a partition representing the invisible symmetry of the output tensor, then we can reduce the number of assignment operations by $\frac{1}{n!}$.
@@ 4.1. Symmetrization
Let $S_{T}|E$ represent the unique symmetry group, which given an equivalence group $E$, consists of $S_{T}|E=\{\pi\in S_{n}\mid\forall i,j\in\{1,2,...,n\},\text{ if }i,j\\ \text{ are both in the same subset of }E,\text{ then }\pi(i)<\pi(j)\}.$
@@ 4.2. Optimization
After symmetrizing the kernel such that it accesses only the canonical triangle(s) of the symmetric tensor(s), we shift to applying various transforms to reduce the number of computations performed. These transforms are the building blocks for filtering redundant code. Since these transformations are performed at the level of sparse tensor computation in Finch IR before Finch lowers to Julia and then LLVM IR, we cannot rely on the Julia or LLVM compilers to perform these optimizations.
@@ 4.2.1. Common Tensor Access Elimination
We replace repeated reads of the same element in a tensor with a single constant value. In particular, after normalizing the symmetrized kernel, all accesses to a fully symmetric tensor will be equivalent in each iteration of a loop. For a fully symmetric tensor of order $n$, this will entail reducing memory reads by $\frac{1}{n!}$. Accesses to other tensors might also be repeated and thus, can also be consolidated. This step is also crucial for the Finch compiler because it understands each tensor access as a separate iterator, and multiple redundant accesses would lead to intersecting multiple iterators in the loop.
@@ 4.2.2. Restrict Computation of Output to Canonical Triangle
Identify assignments with equivalent right-hand sides that update symmetric entries of the output tensor (i.e. coordinates with particular indices swapped) in the same conditional block. In this case, replace the symmetric assignments with just one assignment to the canonical coordinate of the output tensor. We also mark the indices across which the output tensor will need to be replicated. After the kernel is computed, the canonical triangle of the output tensor is replicated to the noncanonical triangles.
@@ 4.2.3. Concordize Tensors
Transpose tensors to make the iteration of indices concordant (Ahrens et al., 2023)—in other words, so that the order of the indices with which the tensor is accessed aligns with the loop order of the kernel. If necessary, transpose the tensor and reorder the loops to make iteration concordant. This step is critical for sparse tensors, as we can only iterate over hierarchical sparse formats with a concordant traversal.
@@ 4.2.4. Consolidate Conditional Blocks
Identify conditional blocks containing equivalent assignments and replace them with a single conditional block with an if-condition that is the union of the if-conditions of each of the conditional blocks. This transform improves the readability of the generated kernel and also prevents unnecessary specialization of cases during compilation.
@@ 5.2. Results
All experiments were run on a single core of a 12-core 2-socket Intel Xeon E5-2695 v2 running at 2.40GHz with 128GB of memory. We used v0.6.22 of the Finch library to implement the kernels and executed both the naive and optimized implementation generated by SySTeC. We compare all kernels to the column-major implementations in TACO, and additionally SSYMV to MKLSparse v1.1.0 and MTTKRP to SPLATT. We used Julia v1.10 to run the tests and all timings are the minimum of 10,000 runs or 5s of measurement, whichever happens first.
For the SSYMV, SYPRD, and SSYRK kernels, we evaluated with the matrix benchmark suite used by Vuduc et. al (Vuduc et al., 2002) and downloaded from the SuiteSparse matrix repository. The asymmetric matrices in the suite were symmetrized by summing the transpose (i.e. $A+A^{T}$). To our knowledge, there does not exist a database of symmetric tensors so for the MTTKRP kernels, we generated uniformly distributed symmetric random sparse tensors of varying sizes and sparsities via an Erdős–Rényi distribution. The dense input matrices are also randomly generated. The numerical rank is the number of columns in the dense matrix, where applicable.
In Figures 5-9, we normalize all results to naive Finch; the red line specifies the performance of naive Finch (our baseline) and the purple line the speedup we expect.
@@ 5.2.1. SSYMV
The sparse symmetric matrix vector kernel is given by $y[i]=A[i,j]*x[j]$ where A is symmetric and CSF, and y and x are dense.
@@ 5.2.4. TTM
The tensor times matrix kernel is given by $C[i,j,l]+=A[k,j,l]*B[k,i]$ where A is fully symmetric CSF, and B and C are dense.
@@ 5.2.5. MTTKRP
The assignments for the 3-, 4-, and 5-dimensional matricized tensor times Khatri-Rao product kernels are given below where A is CSF and B and C are dense.
@@ 6. Related Work
Most of the work on symmetric compilers reduces symmetric problems to other kernels. The most notable of these is the Cyclops Tensor Framework (CTF). CTF reduces $N$-dimensional dense symmetric contractions to $N!$ separate triangular contractions, linearizes triangular indices which are preserved in the output, then dynamically loops over the remaining triangular indices and repeatedly calls matrix multiply, saving compute and storage (Solomonik et al., 2013; Solomonik and Hoefler, 2015). However, this approach does not extend to kernels which are not contractions (such as MTTKRP), and cannot benefit from within-kernel reuse of the redundant arguments that are produced by the problem reduction. This approach also requires transposing and reformatting arguments before running the kernel, which may be expensive in comparison to the cost of the kernel. It also only supports one sparse argument at a time (likely due to the complexity of sparse-sparse interactions in the triangular loops). The OpMin system provides an operation optimization process that identifies the optimal ordering of tensors in a tensor contraction, but does not produce the code to compute the contraction (Lai et al., 2012).
Few compilers produce code that directly computes symmetric kernels. Shi et. al. proposes to use an output-oriented loop structure which iterates through the unique inputs needed to compute the result, but the corresponding random access of the symmetric input resulted in poor performance (Shi et al., 2021). STUR symbolically optimizes kernels based on the structure (triangular, banded, symmetric, etc.) of the arguments using a term rewriting approach, but only applies to static sparsity patterns, and is not as specialized to symmetry (Ghorbani et al., 2023). Similarly, Spampinato proposes a polyhedral approach (sBLACs) to generating structured code, but does not address dynamic sparsity or even tensors of dynamic sizes (Spampinato and Püschel, 2016).
@@ Acknowledgements.
Intel and NSF PPoSS Grant CCF-2217064; DARPA PROWESS Award HR0011-23-C-0101; NSF SHF Grant CCF-2107244; DoE PSAAP Center DE-NA0003965

### v2

@@ Abstract.
copyright: ccdoi: 10.1145/3696443.3708919journalyear: 2025isbn: 979-8-4007-1275-3/25/03conference: Proceedings of the 23rd ACM/IEEE International Symposium on Code Generation and Optimization; March 01–05, 2025; Las Vegas, NV, USAbooktitle: Proceedings of the 23rd ACM/IEEE International Symposium on Code Generation and Optimization (CGO ’25), March 01–05, 2025, Las Vegas, NV, USAccs: Mathematics of computing Mathematical softwareccs: Software and its engineering Compilersccs: Computing methodologies Symbolic and algebraic manipulation
@@ 3.2.1. Visible Output Symmetry
For example, let us consider the first block of the symmetrized TTM kernel given in Listing 1 that performs the assignments using coordinates of $A$ in the canonical triangle that are not on a diagonal. We reorder the assignments to make the pattern from output symmetry more obvious in Listing 2. Swapping the second and third indices in the output tensor on the left-hand side lends an equivalent right-hand side for each expression. As depicted in Listing 3, we can exploit the output symmetry by only writing to the canonical triangle of the output tensor (i.e. if we index C as C[i, j, l], then only where j <= l), which reduces the number of computations that are done by a factor of $2$. Then, we can copy the values from the canonical triangles to the other triangles of the output tensor in a separate loop nest (lines 7-9 of Listing 3), thus completing the remaining assignments.
@@ 9 C[i, j, l] = C[i, l, j]
In general, if $n$ indices are in the same part of a partition representing the visible symmetry of the output tensor, then we can reduce the number of operations by a factor of $n!$.
@@ 4 A = A[i, j]
If $n$ indices are in the same part of a partition representing the invisible symmetry of the output tensor, then we can reduce the number of operations by a factor of $n!$.
@@ 4.1. Symmetrization
Let $S_{T}|E$ represent the unique symmetry group, which given an equivalence group $E$, consists of $S_{T}|E=\{\pi\in S_{n}\mid\forall i,j\in\{1,2,...,n\}$, if $i,j$ are both in the same subset of $E$, then $\pi(i)<\pi(j)\}.$
@@ 4.2. Optimization
After symmetrizing the kernel such that it accesses only the canonical triangle(s) of the symmetric tensor(s), we shift to applying various transforms to reduce the number of computations performed. These transforms are the building blocks for filtering redundant code. Several of these transformations may be familiar, but these transformations are performed at the level of sparse tensor computation in Finch IR before Finch lowers to Julia and then LLVM IR. Therefore, we cannot rely on the Julia or LLVM compilers to perform these optimizations. The simplicial lookup tables and diagonal splitting are novel, as is the overall composition of these transforms with the goal of exploiting symmetry. The compiler is summarized in Figure 4.
Figure 4. Symmetric Compiler Flow
@@ 4.2.1. Common Tensor Access Elimination
We replace repeated reads of the same element in a tensor with a single constant value. In particular, after normalizing the symmetrized kernel, all accesses to a fully symmetric tensor will be equivalent in each iteration of a loop. For a fully symmetric tensor of order $n$, this will entail reducing memory reads by a factor of $n!$. Accesses to other tensors might also be repeated and thus, can also be consolidated. This step is crucial for the Finch compiler because it understands each tensor access as a separate iterator, and multiple redundant accesses would lead to intersecting multiple iterators in the loop. Thus, this step must be applied before running Finch.
@@ 4.2.2. Restrict Computation of Output to Canonical Triangle
Identify assignments with equivalent right-hand sides that update symmetric entries of the output tensor (i.e. coordinates with particular indices swapped) in the same conditional block. In this case, replace the symmetric assignments with just one assignment to the canonical coordinate of the output tensor. We also mark the indices across which the output tensor will need to be replicated. After the kernel is computed, the canonical triangle of the output tensor can be replicated to the noncanonical triangles if needed. We chose to do the replication in a separate loop as the main loop may access the same output location multiple times. By keeping these loops separate, we avoid repeating the out-of-order access required to replicate the output tensor.
@@ 4.2.3. Concordize Tensors
Transpose tensors to make the iteration of indices concordant (Ahrens et al., 2023). A program is concordant when the order of indices in each tensor access match the order in which loops are nested around it. If necessary, transpose the tensor and reorder the loops to make iteration concordant. This step is critical for sparse tensors, as we can only iterate over hierarchical sparse formats with a concordant traversal. Concordant traversal is faster for dense tensors as well. Given a particular loop order and set of tensors, we prioritize transposing the tensors with fewer dimensions (and/or modifying the loop order if necessary) over those with more dimensions for efficiency.
@@ 4.2.4. Consolidate Conditional Blocks
Identify conditional blocks containing equivalent assignments and replace them with a single conditional block with an if-condition that is the union of the if-conditions of each of the conditional blocks. This transform improves the readability of the generated kernel and also prevents unnecessary specialization of cases during Finch compilation.
@@ 5.2. Results
All experiments were run on a single core of a 12-core 2-socket Intel Xeon CPU E5-2680 v3 running at 2.50GHz with 128GB of memory. We used v0.6.22 of the Finch library to implement the kernels and executed both the naive and optimized implementation generated by SySTeC. We compare all kernels to the column-major implementations in TACO, and additionally SSYMV to MKLSparse v1.1.0 and MTTKRP to SPLATT. We used Julia v1.10 to run the tests and all timings are the minimum of 10,000 runs or 5s of measurement, whichever happens first. The time to rearrange data before or after each kernel is not included in the timings, including transposition or replicating the output.
Name Dimension Nonzeros
bayer02 13935 63679
bayer10 13436 94926
bcsstk35 30237 1450163
coater2 9540 207308
crystk02 13965 968583
crystk03 24696 1751178
ct20stif 52329 2698463
ex11 16614 1096948
finan512 74752 596992
gemat11 4929 33185
goodwin 7320 324784
lhr10 10672 232633
lnsp3937 3937 25407
memplus 17758 126150
nasasrb 54870 2677324
Name Dimension Nonzeros
olafu 16146 1015156
onetone2 36057 227628
orani678 2529 90185
raefsky3 21200 1488768
raefsky4 19779 1328611
rdist1 4134 94408
rim 22560 1014951
saylr4 3564 22316
sherman3 5005 20033
sherman5 3312 20793
shyy161 76480 329762
venkat01 62424 1717792
vibrobox 12328 342828
wang3 26064 177168
wang4 26068 177196
Table 2. Matrix collection used in the experiments, taken from Vuduc et. al (Vuduc et al., 2002).
For the SSYMV, SYPRD, and SSYRK kernels, we evaluated with the matrix benchmark suite used by Vuduc et. al (Vuduc et al., 2002), shown in Table 2. The asymmetric matrices in the suite were symmetrized by summing the transpose (i.e. $A+A^{T}$). To our knowledge, there does not exist a database of symmetric tensors so for the MTTKRP kernels, we generated uniformly distributed symmetric random sparse tensors of varying sizes and sparsities via an Erdős–Rényi distribution. The dense input matrices are also randomly generated. The numerical rank is the number of columns in the dense matrix, where applicable.
In Figures 6-11, we normalize all results to naive Finch; the red line specifies the performance of naive Finch (our baseline) and the purple line the speedup we expect.
@@ 5.2.1. SSYMV
The sparse symmetric matrix vector kernel is given by $y[i]=A[i,j]*x[j]$ where A is symmetric and CSC, and y and x are dense.
@@
5.2.2. Bellman-Ford Update
The sparse symmetric Bellman-Ford update kernel is given by $y[i]min=A[i,j]+d[j]$ where A is symmetric and CSC, and y and d are dense. Here, $d$ represents shortest path lengths after $k$ steps, $A$ represents the edge distances, and $y$ represents the shortest path lengths after $k+1$ steps. This kernel is identical to SSYMV from a performance perspective, as seen in Figure 7, but is included to show that SySTeC can handle symmetrizing operations beyond $+$ and *.
Bellman Performance
Figure 7. Bellman-Ford Step Performance
@@ 5.2.5. TTM
The tensor times matrix kernel is given by $C[i,j,l]=A[k,j,l]*B[k,i]$ where A is fully symmetric CSF, and B and C are dense.
@@ 5.2.6. MTTKRP
The MTTKRP (matricized tensor times Khatri-Rao product) kernel is used in tensor factorization, namely to compute the CPD (Canonical Polyadic Decomposition) of a tensor (Kolda and Bader, 2009). Typically, the factorization of an $N$-dimensional tensor requires $N$ MTTKRP kernels, each executed on a different transposition of the tensor. When the tensor is symmetric, no transpose is required because all transpositions of the tensor are equivalent. While CPD typically involves separate factor matrices for each dimension, the symmetric CPD problem uses the same factor matrix for all dimensions. This means that the update step is identical in all the modes, and the algorithm is more efficient (Kofidis and Regalia, 2002, Algorithm 2).
The assignments for the 3-, 4-, and 5-dimensional kernels are given below where A is CSF and B and C are dense.
@@ 6. Related Work
Several tensor compilers support sparse tensor operations, but have no support for tensor symmetry, such as TACO (Kjolstad et al., 2017), SparseLNR (Dias et al., 2022), SPARTA (Liu et al., 2021), and MLIR (Bik et al., 2022). We therefore expect these systems would perform similarly to TACO in our evaluation.
@@
Most of the work on symmetric compilers reduces symmetric problems to other kernels. The most notable of these is the Cyclops Tensor Framework (CTF). CTF reduces $N$-dimensional dense symmetric contractions to $N!$ separate triangular contractions, linearizes triangular indices which are preserved in the output, then dynamically loops over the remaining triangular indices and repeatedly calls matrix multiply, saving compute and storage (Solomonik et al., 2013; Solomonik and Hoefler, 2015). However, this approach does not extend to kernels which are not contractions (such as MTTKRP) or do not use $+$ and * (such as min-$+$ semiring multiplication (Buluç et al., 2017)), and cannot benefit from within-kernel reuse of the redundant arguments that are produced by the problem reduction. This approach also requires transposing and reformatting arguments before running the kernel, which may be expensive in comparison to the cost of the kernel. It also only supports one sparse argument at a time (likely due to the complexity of sparse-sparse interactions in the triangular loops). The OpMin system provides an operation optimization process that identifies the optimal ordering of tensors in a tensor contraction, but does not produce the code to compute the contraction (Lai et al., 2012).
Few compilers produce code that directly computes symmetric kernels. Shi et. al. proposes to use an output-oriented loop structure which iterates through the unique inputs needed to compute the result, but the corresponding random access of the symmetric input resulted in poor performance (Shi et al., 2021). STUR symbolically optimizes kernels based on the structure (triangular, banded, symmetric, etc.) of the arguments using a term rewriting approach, but cannot handle dynamic sparsity patterns, and is not as specialized to symmetry (Ghorbani et al., 2023). Similarly, Spampinato proposes a polyhedral approach (sBLACs) to generating structured code, but does not address dynamic sparsity or even tensors of dynamic sizes (Spampinato and Püschel, 2016). The ITensor library has several library routines for symmetric tensors within tensor networks, but these routines do not handle unstructured sparsity either (Fishman et al., 2022).
@@
7. Future Work
Several avenues for future research remain:
  1. (1)
    Generalizing to More Types of Symmetry: We can expand and adapt our methodology to encompass other forms of symmetry like antisymmetry, block symmetry, or cyclic symmetry that commonly arise in physics, mathematics, chemistry, and machine learning.
  2. (2)
    Exploring Parallelization Opportunities: Incorporating parallelization is important for increasing the practicality of our system, and are actively working towards that goal. In distributed memory settings, the block-cyclic distribution employed by Cyclops (Solomonik et al., 2013) would decompose the problem into several smaller, symmetric instances that we could process with SySTeC out of the box. In multicore settings, we plan to tag an outer loop as parallel. Of course, because iteration over the canonical triangle modifies different transpositions of the output at once, we will also need to make atomic updates to the output.
  3. (3)
    Fine-Tuning Symmetric optimizations: Data formats tailored to symmetry that provide more efficient access when iterating over different transpositions of the canonical triangle could significantly improve performance. Symmetry-aware formats could also eliminate or simplify extra post-processing steps like replicating the canonical triangle of a tensor to the noncanonical triangles. Additionally, our implementation of Systec fully optimizes for every symmetry in the kernel, but we could develop approaches to partially optimize for symmetry only when it is beneficial.
@@ 8. Conclusion
9. Data Availability Statement
The data that support the findings of this study are openly available in Zenodo at https://zenodo.org/records/13821280. (Patel et al., 2024).
@@ Acknowledgements.
Intel and NSF PPoSS Grant CCF-2217064; DARPA PROWESS Award HR0011-23-C-0101; NSF SHF Grant CCF-2107244; DoE PSAAP Center DE-NA0003965;
DARPA SBIR HR001123C0139
Appendix A Artifact Appendix
A.1. Abstract
This artifact provides all of the necessary tools to reproduce the program compilation and charts presented in the paper, with the exception of SSYRK (which takes too much time and memory), and MKL, which is too cumbersome to install. The datasets given to TTM and MTTKRP are smaller here than in the paper. We claim that all speedups should be within 1.5 of the expected, with few exceptions. We note that there was a bug in the TACO TTM results in the submitted version which has been fixed here, so the results may not match the submitted version, but do match the camera ready.
A.2. Artifact check-list (meta-information)
  • Algorithm: We present a new compilation algorithm, SySTeC, which accepts pointwise einsums with symmetry-annotated tensors and produces symmetry-optimized Finch tensor programs.
  • Program: We benchmark on the SSYMV, SYPRD, TTM, and MTTKRP Kernels. We compare against naive Finch, TACO, and SPLATT, all of which are included.
  • Compilation: We require Finch v0.6.32 and Julia v1.10.4 to compile.
  • Binary: We provide a Docker image systec-artifact.tar with all dependencies pre-installed, which requires ARM hardware to run as it was built for Apple M2. If different hardware is used, one can build the docker image for that hardware using the provided instructions.
  • Data set: Matrix datasets are downloaded automatically from http://sparse.tamu.edu/.
  • Run-time environment: The artifact requires a Unix System (Mac or Linux).
  • Hardware: We benchmark on an Intel Xeon CPU E5-2680 v3 running at 2.50GHz, and our benchmarks run well on Apple M2.
  • Execution: We assume that TACO and SPLATT will run in single-threaded mode.
  • Metrics: We measure runtime of all methods we test.
  • Output: The numerical output of each kernel is collected and validated against comparison methods in the script itself. The experiments themselves output json files in the same subdirectory they run in. The artifact includes instructions to plot the results.
  • Experiments: We claim that our speedup results should be within 1.5 of the expected, with few exceptions.
  • How much disk space required (approximately)?: 4GB
  • How much time is needed to prepare workflow (approximately)?: The artifact can be set up quickly with Docker, in around an hour. Building by hand also takes less than an hour, but it may take longer if your system needs special build flags.
  • How much time is needed to complete experiments (approximately)?: The experiments take around an hour to run.
  • Publicly available?: yes
  • Code licenses (if publicly available)?: MIT
  • Workflow framework used?: Docker
  • Archived (provide DOI)?: Zenodo doi.org/10.5281/zenodo.13377504
A.3. Description
A.3.1. How delivered
There are multiple sources:
Download from Zenodo: doi.org/10.5281/zenodo.13377504
Disk space required: 4GB
A.3.2. Hardware dependencies
These experiments require at least 24GB of memory and 4GB disk space. A network connection is required to download the matrix datasets.
We benchmark on an Intel Xeon E5-2695 v2 @ 2.40GHz, and our benchmarks run well on Apple M2.
A.3.3. Software dependencies
A Unix (Mac or Linux) system is required to run the artifact. ARM hardware is required to use the pre-built docker image included with the artifact, or one can build their own docker image or build from source.
Dependencies which must be downloaded are documented in Dockerfile:
  • Julia v1.10.4
  • coreutils
  • cmake
  • gcc
  • g++
  • python
  • python3
  • python3-pip
  • python3-venv
  • git
  • libblas-dev
  • liblapack-dev
  • poetry
Julia deps are also collected in a Project.toml file which can be installed with julia setup.jl or just make env.
Python deps are collected in a pyproject.toml file which can be installed with poetry install --no-root or just make env.
  • python = v3.9
  • matplotlib = v3.9.0
  • packaging = v24.0
  • numpy = v1.26.4
A.3.4. Data sets
All matrix datasets are automatically downloaded with a network connection from http://sparse.tamu.edu/
The tensor datasets are randomly generated by the benchmarks themselves.
A.4. Installation
A.4.1. Docker
Installation is greatly simplified if you are using the provided Docker image. First load the image with
docker image load systec-artifact.tar
or build it with
docker build -t systec-artifact .
Then, you can launch a shell in the container named ”evaluator-container” with
docker run -it --name evaluator systec-artifact
    /bin/bash
Be sure to configure Docker to use enough memory (24GB) and disk space (100 GB) for the experiments.
A.4.2. Manual
Once the basic dependencies are installed on your system, you should be able to run make to build all the necessary dependencies.
Should you need to customize the workflow, make deps builds the dependencies, make env initializes Python and Julia Environments, and make kernels builds the benchmarking kernels.
A.5. Experiment workflow
There are three main phases to the experiments:
A.5.1. Run SySTeC to compile the kernels
From the toplevel directory, you can run the SySTeC compiler with
julia run_SySTeC.jl
This will compile all the kernels required for the experiments and output them into the generated/ directory. The script itself can be inspected and modified if you wish to compile different kernels and compare the output code. The SySTeC compiler itself is contained in deps/SySTeC/src/SySTeC.jl.
A.5.2. Run the experiements
You can run the experiments with
sh run_benchmarks.sh
Benchmarks took us a little over an hour to run. The benchmarks may be run individually with the individual commands in the run_benchmarks.sh script. Results will be generated in the toplevel directory. Corresponding reference results used in the publication are available with the filename suffix _reference.json.
A.5.3. Plot the results
You can plot the results with
poetry run python plot_results.py
This will generate plots for whichever experiments you have run and saved results for. The plots will be saved in the charts/ directory. Corresponding reference results used in the publication are available with the filename suffix _reference.json.
A.6. Evaluation and expected result
You should expect to see similar results to the paper, notably SSYMV (Figure 6), SYPRD (Figure 8), TTM (Figure 10), and MTTKRP (Figure 11). We expect that the more dramatic speedups will be seen in the MTTKRP benchmarks, as the compiler is able to exploit the most symmetry in 5-dimensional kernels. Note that we have reduced the size of the tensors in TTM and MTTKRP to keep the runtime and storage manageable, so these results may demonstrate slightly less speedup as more time is spent on diagonal edge cases.
If you are using our provided Docker image or using Docker to run the experiments, you can view the charts by copying them to the host.
docker cp evaluator:symmetry-benchmarks/charts .
A.7. Experiment customization
The evaluator may optionally customize the kernel given to SySTeC by modifying the run_SySTeC.jl script. The size of the tensors in our TTM and MTTKRP experiments can be changed by modifying the scripts directly.
A.8. Methodology
Submission, reviewing and badging methodology: